OpenStudy (anonymous):

let f(x)= 2x^4 - 4x^2 find all c in the interval (-3, 3) such that f'(c)=0 please show all steps!

myininaya (myininaya):

find f' f'(x)=8x^3-8x f'(c)=8c^3-8c not set f'(c)=0 and solve for c 8c^3-8c=0

OpenStudy (anonymous):

\[f'(x)=8x^3-8x=8x(x^2-1)=8x(x-1)(x+1)=0\]This is true when x=0, x=1, x=-1 all of which are in the interval (-3,3)

OpenStudy (anonymous):

\[f(x)=2x^4-4x^2=0\]\[\Leftrightarrow f(x)=2x^2(x^2-2)=0\] \[\Leftrightarrow 2x^2=0 \Leftrightarrow x=0\] OR \[x^2-2=0\]\[\Leftrightarrow x^2=2\]\[\Leftrightarrow x = \pm \sqrt{2}\]

myininaya (myininaya):

it wan't to know when f'=0

myininaya (myininaya):

he wants*

myininaya (myininaya):

not f

OpenStudy (anonymous):

agreed